# DeepWriter: A Fact-Grounded Multimodal Writing Assistant Based On Offline Knowledge Base

### v1

@@ 6 Conclusion
Knowledge Limitations
DeepWriter’s knowledge is inherently limited by the content of the curated corpus. This restriction ensures reliability but may miss recent developments or emerging trends. Future work will explore hybrid approaches that combine the reliability of offline knowledge bases with the timeliness of online information sources, implementing robust fact-checking mechanisms to maintain quality standards.
Temporal Reasoning
As demonstrated in our case studies, the current system lacks sophisticated temporal reasoning capabilities. Future versions will incorporate temporal understanding to handle queries involving time-sensitive information, policy changes, and historical trends. This includes developing temporal annotation schemes and reasoning modules that can track changes over time within the corpus.
Multimodal Complexity
The current multimodal placement algorithm works well for simple visual elements but struggles with complex charts or diagrams that span multiple topics. Future work will develop more sophisticated visual understanding capabilities, including chart parsing, data extraction, and cross-modal reasoning to better integrate complex visual content.
Scalability and Efficiency
While our hierarchical knowledge representation improves retrieval efficiency, scaling to very large corpora remains a challenge. Future work will explore more efficient indexing strategies, distributed processing approaches, and adaptive chunking algorithms to handle corpora with millions of documents.
Personalization and Customization
The current system generates documents with a general professional tone. Future work will develop user preference modeling and style adaptation capabilities to generate documents tailored to specific audiences, writing styles, and expertise levels.
Interactive Refinement
Current DeepWriter operates in a single-pass generation mode. Future work will explore interactive refinement capabilities, allowing users to provide feedback and guide the generation process iteratively, similar to human writing workflows.
These limitations represent natural extensions of our current work and provide clear research directions for advancing the field of AI-assisted document generation toward more sophisticated, reliable, and user-friendly systems.
@@ Appendix A Prompt of DeepWriter
Section Content Generation Prompt You are an expert research writer tasked with generating high-quality content for a specific section of a comprehensive report. Query: {query} Section Title: {section_title} Section Draft: {section_draft} Relevant Documents: {relevant_docs} Content Already Written in Previous Sections: {already_written} Instructions: 1. Generate detailed, well-structured content for the "section_title" section that directly addresses the query 2. Incorporate information from the relevant documents, synthesizing and analyzing the data 3. Ensure continuity with content already written in previous sections 4. Use an academic, professional tone appropriate for a research report 5. Be thorough but concise, focusing on information that is most relevant to the query 6. Avoid repetition of content already covered in previous sections 7. Do not include title in any level just write the content Your content should: Present factual information directly derived from the relevant documents Synthesize and organize information from multiple sources Maintain neutrality when presenting evidence and data
Summarization Prompt You are an expert summarizer. Your task is to create a concise and accurate summary of the following content in relation to a specific query. The summary should: 1. Capture the main points and key information relevant to the query 2. Highlight the relationship between the content and the query, if there is no relationship, return "None" 3. Maintain the original meaning and intent 4. Be clear and coherent 5. Be no more than 30 percent of the original length Query: {query} Content to summarize: {doc} Provide your summary below, focusing on aspects that address the query:
@@ Appendix B Key Subroutines
1:  $\mathcal{T}_{\text{fact}}\leftarrow\emptyset$
2:  $\mathcal{T}_{\text{data}}\leftarrow\emptyset$
3:  $\mathcal{T}_{\text{point}}\leftarrow\emptyset$
@@
15:  $\mathcal{T}_{\text{point}}\leftarrow\text{IdentifyKeyArguments}(Q^{\prime})$
@@
[…]
@@
13:        $_{k}\leftarrow\text{best\_pos}$_pos
@@
[…]
@@
[…]
@@
[…]

### v2

@@ Appendix A Prompt of DeepWriter
Section Content Generation Prompt You are an expert research writer tasked with generating high-quality content for a specific section of a comprehensive report. Query: {query} Section Title: {section_title} Section Draft: {section_draft} Relevant Documents: {relevant_docs} Content Already Written in Previous Sections: {already_written} Instructions: 1. Generate detailed, well-structured content for the ”section_title” section that directly addresses the query 2. Incorporate information from the relevant documents, synthesizing and analyzing the data 3. Ensure continuity with content already written in previous sections 4. Use an academic, professional tone appropriate for a research report 5. Be thorough but concise, focusing on information that is most relevant to the query 6. Avoid repetition of content already covered in previous sections 7. Do not include title in any level just write the content Your content should: Present factual information directly derived from the relevant documents Synthesize and organize information from multiple sources Maintain neutrality when presenting evidence and data
Summarization Prompt You are an expert summarizer. Your task is to create a concise and accurate summary of the following content in relation to a specific query. The summary should: 1. Capture the main points and key information relevant to the query 2. Highlight the relationship between the content and the query, if there is no relationship, return ”None” 3. Maintain the original meaning and intent 4. Be clear and coherent 5. Be no more than 30 percent of the original length Query: {query} Content to summarize: {doc} Provide your summary below, focusing on aspects that address the query:
@@ Appendix B Key Subroutines
[…]
@@
[…]
@@
10:  $\text{best\_pos}\leftarrow\arg\max_{i}\text{Sim}[k][i]$
11:  $\text{constraints}\leftarrow\text{CheckFlowConstraints}(m_{k},\text{best\_pos},P)$
@@
13:   $\text{pos}_{k}\leftarrow\text{best\_pos}$
@@
4:  $\text{score}\leftarrow\text{SemanticSimilarity}(c,r.\text{content})$
5:  if $\text{score}>\text{highest\_score}$ then
6:   $\text{highest\_score}\leftarrow\text{score}$
7:   $\text{best\_match}\leftarrow r$
@@
12:if $\text{citation.level}=\text{'document'}$ then
13:  $\text{citation.reference}\leftarrow\text{best\_match.filename}$
14:else if $\text{citation.level}=\text{'paragraph'}$ then
15:  $\text{citation.reference}\leftarrow\text{best\_match.filename}+\text{':'}+\text{best\_match.paragraph\_id}$
@@
17:  $\text{citation.reference}\leftarrow\text{best\_match.filename}+\text{':'}+\text{best\_match.sentence\_id}$